# Analysing Moral Bias in Finetuned LLMs through Mechanistic Interpretability

### v1

@@ Abstract
Large language models (LLMs) have been shown to internalize human-like biases during finetuning, yet the mechanisms by which these biases manifest remain unclear. In this work, we investigated whether the well-known Knobe effect, a moral bias in intentionality judgements, emerges in finetuned LLMs and whether it can be traced back to specific components of the model. We conducted a Layer-Patching analysis across 3 open-weights LLMs and demonstrated that the bias is not only learned during finetuning but also localized in a specific set of layers. Surprisingly, we found that patching activations from the corresponding pretrained model into just a few critical layers is sufficient to eliminate the effect. Our findings offer new evidence that social biases in LLMs can be interpreted, localized, and mitigated through targeted interventions, without the need for model retraining.
@@ 1 Introduction
An essential component of human moral judgment is the ability to attribute mental states, such as beliefs and intentions, to moral agents (Young et al. 2007). However, the intentionality attribution process is susceptible to cognitive biases. A well-documented example is the Knobe effect, which demonstrates that individuals are more likely to judge an action as intentional when it leads to a negative side effect than when it results in a positive one, even if the agent’s intention remains the same (Knobe 2003).
As Large Language Models (LLMs) become increasingly integrated into decision-making and ethical reasoning tasks (Dubey, Dailisan, and Mahajan 2025; Dillion et al. 2025), it is crucial to investigate whether they also exhibit the Knobe effect. Understanding this can shed light on how closely their moral reasoning aligns with human cognition and values.
In this study, we addressed three key research aims. First (RQ1), we examined whether different LLMs, specifically Llama (Grattafiori et al. 2024), Mistral (Jiang et al. 2023), and Gemma (Team et al. 2024), manifest the Knobe effect in a way comparable to human behavior.
Consistent with prior findings (Itzhak et al. 2024; Raimondi and Gabbrielli 2025), our work confirmed that these models systematically reproduce human biases, particularly after finetuning. Although designed to align models with human preferences, finetuning also causes the internalization of human biases, which can lead to unintended answers.
@@
Finally (RQ3), we explored whether the observed bias can be effectively reduced. To this end, we developed a Layer-Patching algorithm that selectively overwrites internal activations of finetuned models using those from their pretrained version. Our findings indicate that this method significantly reduces the bias without altering model parameters or requiring further training.
@@
This work is structured as follows: Section 2 provides a comparative overview of related work, Section 3 details our experimental setup organized by research question, Sections 4 and 5 present our results, and Section 6 discusses implications for LLM alignment and future work.
@@ 2 Related Work
Research on the Knobe effect (Knobe 2003) revealed that intentionality attribution is shaped by the valence of an action’s outcome, highlighting that such judgments are not purely objective but influenced by evaluative processes. Building on this, a substantial body of psychological and neuroscientific research has examined how impairments in intentionality attribution (due to brain lesions, neurodevelopmental conditions, or psychiatric disorders) are associated with disruptions in social functioning and moral judgment (Baez et al. 2014; Brüne 2005; Georges, Wiener, and Keller 2013; Young et al. 2010; Sarfati, Hardy-Baylé, and C Besche 1997; Starita et al. 2025; Zucchelli et al. 2019).
@@
Beyond behavioral evaluations, recent work has explored structural and mechanistic differences between human cognition and LLMs. Collacciani, Rambelli, and Bolognesi (2024) analyzed divergences in reasoning patterns, while Bonard and Cortal (2024) investigated how affective signals can be integrated into model responses to improve alignment with human expectations. These contributions highlight that human-like moral reasoning in LLMs depends not only on data but also on how internal representations are shaped.
On a more philosophical level, Chalmers (2023) questioned whether LLMs can genuinely possess intentionality or merely simulate it via statistical correlations. This debate underscores ethical and epistemological importance of understanding biases in LLMs, not merely as artifacts of training data but as emergent properties of model architecture.
While much of the prior work focused on detecting biases at the behavioral level, few studies address where such biases are represented within model architectures or how they might be removed. Advances in mechanistic interpretability, such as Layer-Patching (Meng et al. 2022; Conmy et al. 2023; Nanda and Bloom 2022) and model dissection techniques (Zhang and Nanda 2023), provide tools to localize functional subcomponents responsible for specific behavior.
Recent work applied mechanistic interpretability to address bias localization and mitigation in LLMs. Chandna, Bashir, and Sen (2025) investigate bias in LLMs employing zero-out interventions on activations to identify bias-encoding components. However, their approach results in performance degradation across different tasks, highlighting the challenge of selective bias removal without collateral damage. Similarly, Prakash and Roy (2024) propose a feature-based approach for interpreting bias in LLMs. While their work provides valuable insights, their approach requires counterfactual datasets for bias mitigation.
@@ 3 Methodology
We selected three LLMs families for evaluation: Llama-3 (Grattafiori et al. 2024), Mistral (Jiang et al. 2023), and Gemma (Team et al. 2024). These models were chosen for their relevance in contemporary Natural Language Processing (NLP) tasks and their alignment capabilities for instruction-following applications. Moreover, these models vary in architectural design, pretraining corpora, and finetuning procedures, allowing us to assess the generality of the Knobe effect and the role of model-specific bias. For the main study, we used the following models versions: Llama-3.1-8B, Mistral-7B-v0.1, and gemma-2-9b. To understand the impact of model scale, we then conducted an ablation study using Llama-3.2-1B, gemma-2-2b, and gemma-2-27b.
@@
The dataset used in this study is based on moral scenarios introduced by Ngo et al. (2015). Each model was prompted with a structured query aimed at evaluating intentionality attribution in morally charged contexts. The prompt follows this example format:
@@
The primary motivation for using this dataset is to simulate human experimental conditions and enable direct comparison between model and human responses. The scenarios replicate those used in well-established behavioral studies (Knobe 2003; Zucchelli et al. 2019), allowing the authors to assess whether LLMs reproduce human-like moral asymmetries. This supports the broader goal of evaluating cognitive alignment on whether LLMs internalize human moral biases during finetuning.
@@ 3.1 RQ1
To investigate whether LLMs exhibit the Knobe effect, we conducted a controlled study simulating human population responses. Building on previous research that evaluated intentionality attribution in a human population of 283 participants (Ngo et al. 2015), we generated 283 outputs per model to allow meaningful distributional comparisons.
@@
To introduce variability and simulate a distribution similar to human responses, we employed stochastic sampling with a fixed seed of 0 while randomizing the temperature $T$ over 283 generations per model. The parameter $T$ was sampled from a uniform distribution $\mathcal{U}(T_{\text{min}},T_{\text{max}})$ to prevent deterministic completions and allow the model to express a range of plausible interpretations, where $T_{\text{min}}=0.85$ and $T_{\text{max}}=1.15$. This range was chosen to introduce controlled response stochasticity, approximating the natural variability observed in human moral judgement. It balances diversity and coherence, avoiding degeneration known to occur at higher values (Holtzman et al. 2020). Additionally, recent work suggests that even with stochastic sampling, LLMs tend to underrepresent human behavioral diversity (Qiu et al. 2024).
@@
To quantitatively measure the Knobe effect $\Delta_{\text{Knobe}}$ across models for comparison, we define it as the difference in mean intentionality scores between negative $\mu_{\text{neg}}$ and positive $\mu_{\text{pos}}$ conditions.
@@
To assess variability in responses, we also calculate the standard deviation $\sigma$ across the 283 tests when estimating $\mu_{v}$.
@@
Normality of the data was assessed by examining skewness and kurtosis values for each experimental condition. Skewness values were within the acceptable range of $\pm 2$ across all conditions, indicating symmetrical distributions. Kurtosis slightly exceeded this threshold in one condition, suggesting a moderate deviation from normality in terms of tail heaviness. Given the robustness of repeated-measures ANOVA to moderate violations of normality and the absence of extreme outliers, parametric analyses were deemed appropriate (Blanca et al. 2017).
@@
All analyses are performed using JASP Version 0.95.0 (JASP-Team 2025).
@@ 3.2 RQ2
To address whether specific Transformer layers encode the Knobe effect (RQ2), we hypothesized that certain layers may differentially process morally charged inputs, resulting in measurable shifts in their internal activations. This builds on the idea that Transformer layers tend to specialize in particular functions, from syntactic parsing to semantic abstraction (Nanda and Bloom 2022). If moral reasoning follows similar patterns, a subset of layers may disproportionately contribute to moral judgement, potentially encoding bias.
@@
We used the TransformerLens library (Nanda and Bloom 2022) to extract residual stream activations from layer $l$. For each scenario $x$, we denoted the residual stream activation at layer $l$ as:
@@ 4.1 RQ1
To contextualize our findings within established psychological research, we compared in Table 2 the magnitude of the Knobe effect observed in LLMs with human behavioral data collected using similar experimental conditions by Zucchelli et al. (2019). In that study, 22 participants responded to 20 of the 80 scenarios from the same dataset that we used (Ngo et al. 2015). Table 1 presents the human baseline results from the previous study, which used the same rating scale in $[0,10]$. Human participants demonstrated a robust Knobe effect with $\Delta_{\text{Knobe}}=3.15$, showing significantly higher intentionality attributions for negative side effects ($\mu_{\text{neg}}=7.62$) compared to positive ones ($\mu_{\text{pos}}=4.47$).
@@ 4.2 RQ2
The observed concentration of the Knobe effect in mid-to-late Transformer layers also reflects a general property of training dynamics: the Vanishing Gradient problem (Hochreiter 1998). During finetuning, gradients naturally tend to have a greater impact on mid-to-late layers because of the model’s hierarchical architecture where later layers receive stronger gradient updates. This inherent characteristic of gradient flow causes mid-to-late layers to undergo more substantial adaptation during training, making them the primary locations where learned biases, including moral biases, become encoded and persist.
Moreover, Tenney, Das, and Pavlick (2019) show that the top layers of Transformer models focus more on semantic features and task-specific information, further supporting the view that the emergence of biases in higher layers is driven by finetuning.
@@ 4.3 RQ3
Table 3 quantifies the effect across models. In finetuned Llama-3.1-8B and Mistral-7B models, the intentionality attribution gap $\min{\Delta_{\text{patch}}}$ dropped from 1.6 to 0.0 after patching. In Gemma-2-9B, the gap decreased from 3.83 to just 0.03, showing near-complete reduction. These changes confirm that Knobe effect is not only Mechanistically localizable, but also reversible with targeted interventions.
@@ 5 Ablation Study: Impact of Model Scale
This supports the hypothesis that the Knobe effect in LLMs is primarily induced by the finetuning objective and alignment data, rather than model capacity. Furthermore, the strong effect in the 27B model suggests that finetuning not only introduces the bias but also amplifies it in proportion to the model’s ability to internalize abstract moral constructs as explained in Section 4.2. Here, we show that as the number of layers in a model increases, the bias becomes more pronounced, reflecting the internalization of semantic features in higher layers.
Distributional details and KDE plots are provided in the supplementary material, further illustrating the divergence in intentionality judgements across moral valence and model scale.
@@ 6 Conclusions and Future Work
Our analysis provides concrete evidence that social biases in LLMs not only exist but can also be mechanistically localized, challenging the prevailing view that such behaviors are emergent and diffusely represented (Elhage et al. 2022; Olah et al. 2020). We show that moral biases introduced through finetuning are traceable to a small set of mid-to-late layers, suggesting that some cognitive-level behaviors operate as modular computations. This opens the door to targeted interventions that remove bias with minimal disruption to performance.
Using Layer-Patching (Meng et al. 2022; Dar et al. 2023), we demonstrate that such biases can be selectively mitigated post hoc, revealing interpretability not just as a diagnostic tool but as a practical method for model repair. Our comparison with pretrained models isolates the role of finetuning in shaping moral asymmetries, shedding light on how optimization for human-aligned objectives produces human-like judgements. These findings contribute to debates around machine intentionality (Chalmers 2023), even in the absence of consciousness.
Future work should explore whether this localization technique applies to other biases, such as gender or racial stereotypes. It is also essential to develop methods that do not rely on access to the pretrained model and to rigorously evaluate whether interventions compromise generalization or task performance in unrelated domains. By bridging interpretability and fairness, this work advances both the cognitive modeling of LLMs and the tools for their responsible deployment.
7 Acknowledgments
We acknowledge CINECA [^2] https://www.cineca.it for the availability of computing resources and support.
This work was supported by Future AI Research (FAIR) PE01, SPOKE 8 on PERVASIVE AI funded by the National Recovery and Resilience Plan (NRRP).
Supplementary Material
Ablation study analysis

### v2

@@ Abstract
Large language models (LLMs) have been shown to internalize human-like biases during finetuning, yet the mechanisms by which these biases manifest remain unclear. In this work, we investigated whether the well-known Knobe effect, a moral bias in intentionality judgements, emerges in finetuned LLMs and whether it can be traced back to specific components of the model. We conducted a Layer-Patching analysis across 3 open-weights LLMs and demonstrated that the bias is not only learned during finetuning but also localized in a specific set of layers. Surprisingly, we found that patching activations from the corresponding pretrained model into just a critical layer is sufficient to eliminate the effect. Our findings indicate that social biases in LLMs can be interpreted, localized, and mitigated through targeted interventions, without the need for model retraining.
Machine Learning, ICML
@@ 1 Introduction
An essential component of human moral judgment is the ability to attribute mental states, such as beliefs and intentions, to moral agents (Young et al., 2007). However, the intentionality attribution process is susceptible to cognitive biases. A well-documented example is the Knobe effect, which demonstrates that individuals are more likely to judge an action as intentional when it leads to a negative side effect than when it results in a positive one, even if the agent’s intention remains the same (Knobe, 2003).
As Large Language Models (LLMs) become increasingly integrated into decision-making and ethical reasoning tasks (Dubey et al., 2025; Dillion et al., 2025), it is crucial to investigate whether they also exhibit the Knobe effect. Understanding this can shed light on how closely their moral reasoning aligns with human cognition and values.
In this study, we addressed three key research aims. First (RQ1), we examined whether different LLMs, specifically Llama (Grattafiori et al., 2024), Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024), manifest the Knobe effect in a way comparable to human behavior.
Consistent with prior findings (Itzhak et al., 2024; Raimondi & Gabbrielli, 2025), our work confirmed that these models systematically reproduce human biases, particularly after finetuning. Although designed to align models with human preferences, finetuning also causes the internalization of human biases, which can lead to unintended answers.
@@
Finally (RQ3), we explored whether the observed bias can be effectively reduced. To this end, we developed a Layer-Patching algorithm that selectively overwrites internal activations of finetuned models using those from their pretrained version. Our findings indicate that this method reduces the bias without altering model parameters or further training.
@@ 2 Related Work
Research on the Knobe effect (Knobe, 2003) revealed that intentionality attribution is shaped by the valence of an action’s outcome, highlighting that such judgments are not purely objective but influenced by evaluative processes. Building on this, substantial psychological and neuroscientific research has linked impairments in intentionality attribution (due to brain lesions, neurodevelopmental conditions, or psychiatric disorders) to disruptions in social functioning and moral judgment (Baez et al., 2014; Brüne, 2005; Georges et al., 2013; Young et al., 2010; Sarfati et al., 1997; Starita et al., 2025; Zucchelli et al., 2019).
@@
Beyond behavioral evaluations, recent work has explored structural and mechanistic differences between human cognition and LLMs. Collacciani et al. (2024) analyzed divergences in reasoning patterns, while Bonard & Cortal (2024) investigated how affective signals can be integrated into model responses to improve alignment with human expectations. These contributions highlight that human-like moral reasoning in LLMs depends not only on data but also on how internal representations are shaped.
On a more philosophical level, Chalmers (2023) questioned whether LLMs can possess intentionality or simulate it via statistical correlations. This debate underscores ethical and epistemological importance of understanding biases in LLMs, not merely as artifacts of training data but as emergent properties of model architecture.
While much of the prior work focused on detecting biases at the behavioral level, few studies address where such biases are represented within model architectures or how they might be removed. Advances in mechanistic interpretability, such as Layer-Patching (Meng et al., 2022; Conmy et al., 2023; Nanda & Bloom, 2022) and model dissection techniques (Zhang & Nanda, 2023), provide tools to localize functional subcomponents responsible for specific behavior.
Recent work applied mechanistic interpretability to address bias localization and mitigation in LLMs. Chandna et al. (2025) investigate bias in LLMs using zero-out interventions on activations to identify bias-encoding components. However, their approach results in performance degradation across different tasks, highlighting the challenge of selective bias removal without collateral damage. Similarly, Prakash & Roy (2024) propose a feature-based approach that, while insightful, requires counterfactual datasets for bias mitigation.
@@ 3 Methodology
We selected three LLMs families for evaluation: Llama-3 (Grattafiori et al., 2024), Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024). These models were chosen for their relevance in contemporary Natural Language Processing (NLP) tasks and their alignment capabilities for instruction-following applications. Moreover, these models vary in architectural design, pretraining corpora, and finetuning procedures, allowing us to assess the generality of the Knobe effect and the role of model-specific bias. In main study, we used Llama-3.1-8B, Mistral-7B-v0.1, and gemma-2-9b. To understand the impact of model scale, we then conducted an ablation study using Llama-3.2-1B, gemma-2-2b, and gemma-2-27b.
@@
We used data based on moral scenarios introduced by Ngo et al. (2015). Each model was prompted with a structured query aimed at evaluating intentionality attribution in morally charged contexts following this format:
@@
The primary motivation for using this dataset is to simulate human experimental conditions and enable direct comparison between model and human responses. The scenarios replicate those used in well-established behavioral studies (Knobe, 2003; Zucchelli et al., 2019), allowing the authors to assess whether LLMs reproduce human-like moral asymmetries. This supports the broader goal of evaluating cognitive alignment on whether LLMs internalize human moral biases during finetuning.
@@ 3.1 RQ1
To investigate whether LLMs exhibit the Knobe effect, we conducted a controlled study simulating human population responses. Building on previous research that evaluated intentionality attribution in a human population of 283 participants (Ngo et al., 2015), we generated 283 outputs per model to allow meaningful distributional comparisons.
@@
To introduce variability and simulate a distribution similar to human responses, we employed stochastic sampling with a fixed seed of 0 while randomizing the temperature $T$ over 283 generations per model. The parameter $T$ was sampled from a uniform distribution $\mathcal{U}(T_{\text{min}},T_{\text{max}})$ to prevent deterministic completions and allow the model to express a range of plausible interpretations, where $T_{\text{min}}=0.85$ and $T_{\text{max}}=1.15$. This range was chosen to introduce controlled response stochasticity, approximating the natural variability observed in human moral judgement. It balances diversity and coherence, avoiding degeneration known to occur at higher values (Holtzman et al., 2020). Additionally, recent work suggests that even with stochastic sampling, LLMs tend to underrepresent human behavioral diversity (Qiu et al., 2024).
@@
To quantitatively measure the Knobe effect $\Delta_{\text{Knobe}}$ across models, we define it as the difference in mean intentionality scores between negative $\mu_{\text{neg}}$ and positive $\mu_{\text{pos}}$ conditions.
@@
To assess variability in responses, we also calculate the standard deviation $\sigma$ across the 283 tests for $\mu_{v}$.
@@
Normality of the data was assessed by examining skewness and kurtosis values for each experimental condition. Skewness values were within the acceptable range of $\pm 2$ across all conditions, indicating symmetrical distributions. Kurtosis slightly exceeded this threshold in one condition, suggesting a moderate deviation from normality in terms of tail heaviness. Given the robustness of repeated-measures ANOVA to moderate violations of normality and the absence of extreme outliers, parametric analyses were deemed appropriate (Blanca et al., 2017).
@@
All analyses are performed using JASP Version 0.95.0 (JASP-Team, 2025).
@@ 3.2 RQ2
To address whether specific Transformer layers encode the Knobe effect (RQ2), we hypothesized that certain layers may differentially process morally charged inputs, resulting in measurable shifts in their internal activations. This builds on the idea that Transformer layers tend to specialize in particular functions, from syntactic parsing to semantic abstraction (Nanda & Bloom, 2022). If moral reasoning follows similar patterns, a subset of layers may disproportionately contribute to moral judgement, potentially encoding bias.
@@
We used the TransformerLens library (Nanda & Bloom, 2022) to extract residual stream activations from layer $l$. For each scenario $x$, we denoted the residual stream activation at layer $l$ as:
@@ 4.1 RQ1
To contextualize our findings within established psychological research, we compared in Table 2 the magnitude of the Knobe effect observed in LLMs with human behavioral data collected using similar experimental conditions by Zucchelli et al. (2019). In that study, 22 participants responded to 20 of the 80 scenarios from the same dataset that we used (Ngo et al., 2015). Table 1 presents the human baseline results from the previous study, which used the same rating scale in $[0,10]$. Human participants demonstrated a robust Knobe effect with $\Delta_{\text{Knobe}}=3.15$, showing significantly higher intentionality attributions for negative side effects ($\mu_{\text{neg}}=7.62$) compared to positive ones ($\mu_{\text{pos}}=4.47$).
@@ 4.2 RQ2
The observed concentration of the Knobe effect in mid-to-late layers also reflects a general property of Transformer: Tenney et al. (2019) show that the top layers of Transformer models focus more on semantic features and task-specific information, further supporting the view that the emergence of biases in higher layers is driven by finetuning.
@@ 4.3 RQ3
Table 3 quantifies the effect across models. In finetuned Llama-3.1-8B and Mistral-7B models, the intentionality attribution gap $\min{\Delta_{\text{patch}}}$ dropped from 1.6 to 0.0 after patching. In Gemma-2-9B, the gap decreased from 3.83 to just 0.03, showing near-complete reduction. These changes confirm that Knobe effect can be both mechanistically localized and reversed through targeted interventions.
Model Type ARC-Easy HellaSwag MMLU TruthfulQA Model $\Delta$
Llama-3.1-8B Pretrained 0.8152 (+0.0063) 0.5998 (-0.0107) 0.6339 (+0.0096) 0.4514 (+0.0827) +0.0220
Finetuned 0.8190 (+0.0025) 0.5915 (-0.0024) 0.6809 (-0.0374) 0.5406 (-0.0065) -0.0110
Patched 0.8216 0.5891 0.6435 0.5341
Mistral-7B-v0.1 Pretrained 0.8081 (-0.0101) 0.6131 (-0.0173) 0.5949 (-0.0022) 0.4261 (+0.0351) +0.0014
Finetuned 0.8005 (-0.0025) 0.5629 (+0.0329) 0.5335 (+0.0593) 0.5594 (-0.0983) -0.0021
Patched 0.7980 0.5958 0.5927 0.4611
Gemma-2-9B Pretrained 0.8725 (+0.0046) 0.6104 (-0.0212) 0.6904 (+0.0062) 0.4544 (+0.0993) +0.0222
Finetuned 0.8590 (+0.0181) 0.5961 (-0.0069) 0.7193 (-0.0228) 0.6018 (-0.0481) -0.0149
Patched 0.8771 0.5892 0.6966 0.5538
Table 4: Benchmark accuracy of models. The patched models demonstrate strong utility preservation, showing minimal regression on general tasks after bias removal when compared to the finetuned versions. Parenthetical values indicate each model’s accuracy change relative to the patched version. Model $\Delta$ is the average difference over all the benchmarks.
To ensure that Layer-Patching does not compromise general model capabilities, we evaluated performance across four standard benchmarks: ARC-Easy, HellaSwag, MMLU, and TruthfulQA. As shown in Table 4, the intervention avoids catastrophic degradation. All the models maintain performance levels largely comparable to the finetuned baselines, with only minor regressions around 1%. These results confirm that Layer-Patching acts as a surgical intervention, effectively neutralizing moral bias without sacrificing the model’s broader utility.
@@ 5 Ablation Study: Impact of Model Scale
This supports the hypothesis that the Knobe effect in LLMs is primarily induced by the finetuning objective and alignment data, rather than model capacity. The strong effect in the 27B model suggests finetuning amplifies the bias in proportion to the model’s ability to internalize abstract moral constructs, reflecting the increasing internalization of semantic features in higher layers (Section 4.2).
Distributional details are provided in Appendix A, further illustrating the divergence in intentionality judgements across moral valence and model scale.
@@ 6 Conclusions and Future Work
Our analysis provides concrete evidence that social biases in LLMs not only exist but can also be mechanistically localized, challenging the prevailing view that such behaviors are emergent and diffusely represented (Elhage et al., 2021; Olah et al., 2020). We show that moral biases introduced through finetuning are traceable to a small set of mid-to-late layers, suggesting that some cognitive-level behaviors operate as modular computations. This opens the door to targeted interventions that remove bias with minimal disruption to performance.
Using Layer-Patching (Meng et al., 2022; Dar et al., 2023), we demonstrate that such biases can be selectively mitigated post hoc, revealing interpretability not just as a diagnostic tool but as a practical method for model repair. Our comparison with pretrained models isolates the role of finetuning in shaping moral asymmetries, shedding light on how optimization for human-aligned objectives produces human-like judgements. These findings contribute to debates around machine intentionality (Chalmers, 2023), even in the absence of consciousness.
Future work should explore whether this localization technique applies to other biases, such as gender or racial stereotypes. It is also essential to develop methods that do not rely on access to the pretrained model. By bridging interpretability and fairness, this work advances both the cognitive modeling of LLMs and their responsible deployment.
Appendix A Appendix
A.1 Ablation study analysis